Grade 11Algebra

The solution of the equation log base7 log base5(√x+5 +√x)=0
X=??
(a)4
(b)2
(c)3
(d)0

Aryan Nehra
8 Years agoGrade 11

1 Answer

akhil ojha
ApprovedApproved Tutor Answer8 Years ago
log7(log5(x1/2+5+x1/2))=0 this is our  question as we know that 
if logxy=0
then y should be 1
 
so
log5(2x1/2+5)=1
and we also know that logxy=1
then y should be x
2x1/2+5=5
2x1/2=0
x=0